The heater element of a particular 120-V toaster is a 8.9-m length of nichrome wire, whose diameter is 0.86 mm. The resistivity of nichrome at the operating temperature of the toaster is 1.3
Answer #1
720 W


R = rho l /A

R = 1.3 × 10-6 x 8.9 / pi x .43^2 x 10^-6 = 20 ohm

power = V^2 /R = 120 x120 /20 = 720 W